A closed system consisting of 1kg gaseous co² undergoes a reversible process at constant pressure causing a decrease of 30kj in internal energy
Determine the work done during the process
Take CP=840j/kg°c and CV=600j/kg°c
1
Expert's answer
2020-06-16T09:27:49-0400

As per the given question,

m=1 kg

Pressure is constant.

Change in internal energy of the gas =-30kJ

Work one during the process=?

CP=840j/kg°c

CV=600j/kg°c

We know that "\Delta U= mC_v(T_1 -T_2)"

Now substituting the values,

"\Rightarrow -30000J=1\times 600(T_1 -T_2)"

"\Rightarrow (T_1 -T_2)= 50^\circ C"

The supplied heat or rejected heat,

"\Rightarrow Q=m c_p(T_2-T_1)"

"=840\times -50"

"=-42kJ"

We know that,

"Q=\Delta U+W"

"\Rightarrow W=Q-\Delta U"

"\Rightarrow W=-42kJ+30=-12kJ"